### v2

# Climbing over the potential barrier during inflation via null energy condition violation

Abstract
The violation of the null energy condition (NEC) may play a crucial role in enabling a scalar field to climb over high potential barriers, potentially significant in the very early universe. We propose a single-field model where the universe sequentially undergoes a first stage of slow-roll inflation, NEC violation, and a second stage of slow-roll inflation. Through the NEC violation, the scalar field climbs over high potential barriers, leaving unique characteristics on the primordial gravitational wave power spectrum, including a blue-tilted nature in the middle-frequency range and diminishing oscillation amplitudes at higher frequencies. Additionally, the power spectrum exhibits nearly scale-invariant behavior on both large and small scales.
pacs:
98.80.-k, 98.80.Cq, 04.50.Kd
I Introduction
The detection of gravitational waves (GWs) originating from binary pulsar systems Detweiler:1979wn and binary black hole mergers LIGOScientific:2016aoc has opened a new avenue for exploring gravity and the universe. Recently, several collaborations within the pulsar timing array (PTA) community, including NANOGrav NANOGrav:2023gor ; NANOGrav:2023hvm ; NANOGrav:2020bcs , EPTA EPTA:2023fyk , PPTA Reardon:2023gzh and CPTA Xu:2023wog , have announced compelling evidence of a signal consistent with stochastic GW background at the frequency $f\simeq 1~{}\textnormal{yr}^{-1}$.
Inflation, as the most popular paradigm of the early universe, predicts the existence of a primordial GW background Grishchuk:1974ny ; Starobinsky:1979ty ; Rubakov:1982df across a wide frequency range ($10^{-18}-10^{10}$ Hz), yet to be confirmed by observations Guth:1980zm ; Linde:1981mu ; Albrecht:1982wi ; Starobinsky:1980te . It is interesting to ask whether the signals recently observed by PTA could originate from primordial GWs, see, e.g., Vagnozzi:2023lwo ; Jiang:2023gfe ; Zhu:2023lbf ; Ye:2023tpz ; Frosina:2023nxu ; Ellis:2023oxs (see also Huang:2023chx ; Du:2023qvj ; Wu:2023hsa ; Xiao:2023dbb ; Zhang:2023lzt ; Ye:2023xyr ; King:2023ayw ; Huang:2023zvs ; Lozanov:2023rcd ; Maji:2023fhv ; Kawasaki:2023rfx ; He:2023ado ; Choudhury:2023fwk ; Choudhury:2023fjs ; Jiang:2024dxj ; DeAngelis:2024xtr ). The standard slow-roll inflation predicts a nearly scale-invariant power spectrum of primordial GWs, rigorously constrained by Planck’s data, which yields a tensor-to-scalar ratio $r_{0.002}<0.035$ within the observational window of the cosmic microwave background (CMB) BICEP:2021xfz . However, PTA data indicates a highly blue-tilted GW power spectrum in the nanohertz frequency band, with a spectral index of $n_{T}=1.8\pm 0.3$ NANOGrav:2023hvm ; Vagnozzi:2023lwo .
Attributing the observed signal in the PTA band to primordial GWs necessitates physics beyond standard slow-roll inflation, see, e.g., Jiang:2023gfe ; Zhu:2023lbf ; Ye:2023tpz ; Borah:2023sbc ; Ben-Dayan:2023lwd ; Oikonomou:2023qfz ; Datta:2023xpr ; Datta:2023vbs ; Bhattacharya:2023ysp ; Choudhury:2023hfm ; Choudhury:2023kdb ; Choudhury:2024one ; Sharma:2024whg , for recent studies. This would require the amplitude of the primordial GW power spectrum $P_{T}$ to reach a maximum of $P_{T}\simeq 10^{-3}$ at PTA scales (see, e.g., Ye:2023tpz ), where $P_{T}$ is related to the present GW energy density spectrum $\Omega_{\rm{GW}}$ by $P_{T}\simeq 10^{6}\Omega_{\rm{GW}}$. There have been numerous studies exploring the generation of a blue-tilted GW spectrum by introducing new physics beyond conventional slow-roll inflation, e.g., Piao:2004tq ; Baldi:2005gk ; Piao:2006jz ; Kobayashi:2010cm ; Kobayashi:2011nu ; Dudas:2010gi ; Mourad:2017rrl ; Endlich:2012pz ; Cai:2014uka ; Gong:2014qga ; Cannone:2014uqa ; Wang:2014kqa ; Kuroyanagi:2014nba ; Cai:2015yza ; Cai:2016ldn ; Wang:2016tbj ; Fujita:2018ehq ; Kuroyanagi:2020sfw ; Akama:2020jko ; Akama:2023jsb ; Giare:2020plo ; Datta:2022tab ; Giare:2022wxq ; Oikonomou:2023bli ; Choudhury:2023kam ; Oikonomou:2024aww . An intriguing direction is to introduce a violation of the null energy condition (NEC) during inflation.
Realizing a stable NEC violation is a challenge due to the presence of ghost or gradient instabilities in the primordial perturbations associated with the NEC violation Rubakov:2014jja ; Libanov:2016kfc ; Kobayashi:2016xpl ; Easson:2011zy ; Qiu:2015nha ; Ijjas:2016tpn ; Ijjas:2016vtq ; Dobre:2017pnt ; deCesare:2019pqj ; deCesare:2021wmk . It is first explicitly demonstrated in the framework of effective field theory that a fully stable444“Fully stable” implies that both ghost and gradient instabilities are eliminated throughout the entire history of the universe. NEC violation can be realized in “beyond Horndeski” theories Cai:2016thi ; Creminelli:2016zwa ; Cai:2017tku ; Cai:2017dyi ; Kolevatov:2017voe , see also Cai:2017dxl ; Cai:2017pga ; Mironov:2018oec ; Qiu:2018nle ; Ye:2019frg ; Ye:2019sth ; Mironov:2019qjt ; Akama:2019qeh ; Mironov:2019mye ; Ilyas:2020qja ; Ilyas:2020zcb ; Zhu:2021whu ; Zhu:2021ggm ; Mironov:2022ffa ; Cai:2022ori ; Panda:2024iqu for later developments. On this basis, it has been demonstrated that NEC violation during inflation can have significant observable effects, including: 1) an enhanced and blue-tilted power spectrum of primordial GWs with distinct features Cai:2020qpu ; Cai:2022nqv , which could be detected by PTA Ye:2023tpz , Advanced LIGO and Advanced Virgo Chen:2024mwg , and Taiji Chen:2024jca , 2) the generation of primordial black holes and scalar-induced GWs Cai:2023uhc , 3) an amplified parity-violation effect in primordial GWs Cai:2022lec ; Zhu:2023lhv .
In this paper, we investigate a new model where NEC violation occurs during slow-roll inflation, enabling the inflaton to climb over a high potential barrier before transitioning back to slow-roll inflation. Unlike the model in Cai:2020qpu , in this model, the two stages of inflation before and after NEC violation have similar energy scales. We will investigate the observable effects of NEC violation in this model on the primordial GW background through analytical and numerical calculations.
II The model
In our single scalar field model, the universe originates from a slow-roll inflation. As the inflaton rolls slowly down the potential, it encounters a high potential barrier (see Fig. 3 for a sketch of the potential). Typically, a slowly rolling inflaton cannot climb over this high barrier. However, through the mechanism of NEC violation, the inflaton naturally violates the slow-roll conditions for a certain duration, allowing it to climb over the barrier and continue the slow-roll inflation, as illustrated by Fig. 1. Such an evolution will lead to interesting observable effects on the primordial perturbations.
Figure 1: In the first panel, inflaton can climb over the barrier which are not too high. But in the other two panels, inflaton requires the NEC violation to climb over the barrier and continue the slow-roll inflation.
Following Cai:2020qpu , we use the action
$S=\int d^{4}x\sqrt{-g}\Big{[}{M_{p}^{2}\over 2}R-M_{p}^{2}g_{1}(\phi)X/2+g_{2}(\phi)X^{2}/4-M_{p}^{4}V(\phi)+L_{\delta g^{00}R^{(3)}}\Big{]}\,,$ (1)
where $\phi$ is a dimensionless scalar field, $X=\nabla_{\mu}\phi\nabla^{\mu}\phi$, the EFT operator $L_{\delta g^{00}R^{(3)}}$ is adopted to thoroughly eliminate the ghost and gradient instabilities of the primordial scalar perturbations, see, e.g., Cai:2017dyi for details. According to Cai:2016thi ; Cai:2017dyi , $L_{\delta g^{00}R^{(3)}}$ does not affect the background equations at all, nor does it contribute to the quadratic action of the tensor sector. It only modifies the squared sound speed of the scalar perturbations. For simplicity, we will focus on the primordial GWs in this paper.
The background equations can be obtained as
$3H^{2}M_{p}^{2}$ $=$ $\frac{M_{p}^{2}}{2}g_{1}\dot{\phi}^{2}+\frac{3}{4}g_{2}\dot{\phi}^{4}+M_{p}^{4}V\,,$ (2)
$\dot{H}M_{p}^{2}$ $=$ $-\frac{M_{p}^{2}}{2}g_{1}\dot{\phi}^{2}-\frac{1}{2}g_{2}\dot{\phi}^{4}\,,$ (3)
00 0 $=$ $\left(g_{1}+\frac{3g_{2}\dot{\phi}^{2}}{M_{p}^{2}}\right)\ddot{\phi}+3g_{1}H\dot{\phi}+\frac{1}{2}g_{1,\phi}\dot{\phi}^{2}+\frac{3g_{2}H\dot{\phi}^{3}}{M_{p}^{2}}+\frac{3g_{2,\phi}\dot{\phi}^{4}}{4M_{p}^{2}}+M_{p}^{2}V_{,\phi}\,,$ (4)
where “${}_{,\phi}\equiv d/d\phi$”, a dot denotes the derivative with respect to the comic time $t$. In the background equations (2) to (4), only two of them are independent.
(a)  $g_{1}(\phi)$
(b)  $g_{2}(\phi)$
Figure 2: The functions $g_{1}(\phi)$ and $g_{2}(\phi)$, where we have set $q_{1}=10$, $q_{2}=6$, $q_{3}=10$, $\phi_{0}=3.2$, $\phi_{3}=-4.38$, $f_{1}=1$, $f_{2}=40$. A sufficiently large positive $g_{2}$ that satisfies $g_{2}{\dot{\phi}}^{4}>M_{\rm P}^{2}\dot{H}$ can prevent ghost instabilities.
Figure 3: A plot illustrating the potential $V(\phi)$ with parameters set as follows: $\phi_{1}=2$, $\phi_{2}=-4$, $\phi_{4}=20$, $\phi_{5}=29$, $q_{2}=6$, $q_{4}=4$, $\lambda=0.01$, $\sigma=23$, and $m=-4.5\times 10^{-6}$. The height of the potential barrier significantly exceeds that of both sides where slow-roll inflation occurs, i.e., $V_{\mathrm{top}}\gg V_{\mathrm{inf1}}$ and $V_{\mathrm{inf2}}$. As the scalar field rolls along the potential from left to right, the universe sequentially experiences the first stage of slow-roll inflation, NEC violation, and the second stage of slow-roll inflation.
To achieve the aforementioned background evolution via NEC violation, we set these $\phi$-dependent functions $g_{1}$, $g_{2}$ and the potential $V$ in action (1) as
$g_{1}(\phi)$ $=$ ${1\over 1+e^{q_{2}(\phi-\phi_{3})}}-{f_{1}e^{2\phi}\over 1+f_{1}e^{2\phi}}+{2\over 1+e^{-q_{1}(\phi-\phi_{0})}}\,,$ (5)
$g_{2}(\phi)$ $=$ ${f_{2}\over 1+e^{-q_{2}(\phi-\phi_{3})}}{1\over 1+e^{q_{3}(\phi-\phi_{0})}}\,,$ (6)
$\begin{split}V(\phi)=&{1\over 2}m^{2}(\phi-\phi_{4})^{2}{1\over 1+e^{q_{2}(\phi-\phi_{2})}}+\lambda\left[1-\frac{(\phi-\phi_{1})^{2}}{\sigma^{2}}\right]^{2}{e^{-q_{4}(\phi-\phi_{1})}\over{(1+e^{-q_{4}(\phi-\phi_{1})})}^{2}}\\ &+{1\over 2}m^{2}(\phi-\phi_{5})^{2}{1\over 1+e^{-q_{4}(\phi-\phi_{1})}}\,,\end{split}$ (7)
where $\lambda$, $m$, $f_{1,2}$, $q_{1,2,3,4}$ and $\phi_{0,1,2,3,4,5}$ are dimensionless constants. We plot $g_{1}(\phi)$, $g_{2}(\phi)$ and $V(\phi)$ for appropriate values of these parameters in Figs. 2 and 3. As the scalar field rolls along the potential from left to right, the universe sequentially experiences the first stage of slow-roll inflation, NEC violation, and the second stage of slow-roll inflation. These two slow-roll stages are labeled as “inf1” and “inf2” respectively.
The parameters introduced above should satisfy $\phi_{3}<\phi_{2}<0<\phi_{1}<\phi_{0}$. As can be seen in Figs. 2 and 3, when $\phi$ is much smaller than $\phi_{3}$, $g_{1}=1$ and $g_{2}=0$, rendering the potential responsible for the first stage of slow-roll inflation. In this regime, the first term of the potential predominates. Conversely, as $\phi$ greatly exceeds $\phi_{0}$, $g_{1}=1$ and $g_{2}=0$, and the potential is responsible for the second stage of slow-roll inflation, with the third term in the potential becoming dominant. The exponential term in each expression serves to modulate the dominance of various terms across different regions of the potential.
From Fig. 3, it can be observed that the height of the potential barrier significantly exceeds that of both sides where slow-roll inflation occurs. For a slowly rolling canonical scalar field, such a high barrier is insurmountable. However, with the action described in Eq. (1) and via the NEC violation mechanism, the scalar field can naturally disrupt the slow-roll conditions and rapidly climb over the barrier.
The drop of the potential before the hill is significant for the $\phi$ field to climb to the top of the potential. In our model, the equation of motion for $\phi$ is given by Eq. (4). The coefficient in front of $\ddot{\phi}$ is proportional to $Q_{s}H^{2}/\dot{\phi}^{2}$, which must remain positive throughout to avoid ghost instability. To allow the $\phi$ field to climb to the top of the potential without requiring a non-zero initial kinetic energy, we set the model parameters so that $g_{1}<0$ around the NEC-violating phase. This results in the term $3g_{1}H\dot{\phi}$ acting as an anti-friction force, thereby accelerating the $\phi$ field. Therefore, both the magnitude of $\dot{\phi}$ at the beginning of the climb and the coefficient in front of $\ddot{\phi}$ will influence the $\phi$ field’s ability to climb over the hill. This highlights the importance of the drop of the potential before the hill. Otherwise, if $\dot{\phi}$ is too small when the field starts climbing, we might need to finely tune the model parameters to ensure that the $\phi$ field reaches the top of the potential.
We numerically solve the background evolutions and display the evolutions of $\phi$, $\dot{\phi}$, $H$ and $\epsilon\equiv-\dot{H}/H^{2}$ in Figs. 4 and 5. From these figures, we can see that the universe sequentially experiences the first stage of slow-roll inflation, NEC violation, and the second stage of slow-roll inflation. This background evolution is similar to that in the scenario described in Cai:2020qpu . The main difference from the model in Cai:2020qpu lies in the fact that in this model, the energy scales of the first and second stages of slow-roll inflation are closer to each other.
Figure 4: The evolutions of $\phi(t)$ (left panel) and $\dot{\phi}$ (right panel), where $\phi(t_{\rm ini})=-6$, $\dot{\phi}(t_{\rm ini})=0$, $t_{\rm ini}=0$, $\phi_{0}=3.2$, $\phi_{1}=2$, $\phi_{2}=-4$, $\phi_{3}=-4.38$, $q_{1}=10$, $q_{2}=6$, $q_{3}=10$, $q_{4}=4$, $f_{1}=1$, $f_{2}=40$, $\lambda=0.01$, $\sigma=23$ and $m=-4.5\times 10^{-6}$.
Figure 5: The evolutions of a hundredfold of $H$ (blue curve) and $\epsilon=-\dot{H}/H^{2}$ (red curve). We have $0<\epsilon\ll 1$ during the first and second stages of slow-roll inflation, and $\epsilon\ll-1$ during the NEC-violating stage. Note that when $\dot{\phi}/H$ reaches its peak, $H$ is still far from its maximum.
III The power spectrum of Primordial GWs
In this section, we calculate the power spectrum of primordial GWs for our model. The quadratic actions of tensor perturbation mode $\gamma_{ij}$ can be given as
$S^{(2)}_{\gamma}={M_{p}^{2}\over 8}\int d^{4}xa^{3}\left[\dot{\gamma}_{ij}^{2}-{(\partial_{k}\gamma_{ij})^{2}\over a^{2}}\right]\,.$ (8)
In the momentum space,
$\gamma_{ij}(\tau,\mathbf{x})=\int\frac{d^{3}k}{(2\pi)^{3}}e^{-i\mathbf{k}\cdot\mathbf{x}}\sum_{\lambda=+,\times}\hat{\gamma}_{\lambda}(\tau,\mathbf{k})\epsilon^{(\lambda)}_{ij}(\mathbf{k}),$ (9)
where $\hat{\gamma}_{\lambda}(\tau,\mathbf{k})=\gamma_{\lambda}(\tau,k)a_{\lambda}(\mathbf{k})+\gamma_{\lambda}^{*}(\tau,-k)a_{\lambda}^{{\dagger}}(-\mathbf{k})$, $\epsilon_{ij}^{(\lambda)}(\mathbf{k})$ satisfies $k_{j}\epsilon_{ij}^{(\lambda)}(\mathbf{k})=0$, $\epsilon_{ii}^{(\lambda)}(\mathbf{k})=0$, $\epsilon_{ij}^{(\lambda)}(\mathbf{k})\epsilon_{ij}^{*(\lambda^{\prime})}(\mathbf{k})=\delta_{\lambda\lambda^{\prime}}$ and $\epsilon_{ij}^{*(\lambda)}(\mathbf{k})=\epsilon_{ij}^{(\lambda)}(-\mathbf{k})$; $a_{\lambda}(\mathbf{k})$ and $a^{{\dagger}}_{\lambda}(\mathbf{k}^{\prime})$ satisfy $[a_{\lambda}(\mathbf{k}),a_{\lambda^{\prime}}^{{\dagger}}(\mathbf{k}^{\prime})]=\delta_{\lambda\lambda^{\prime}}\delta^{(3)}(\mathbf{k}-\mathbf{k}^{\prime})$.
Using the Eq. (8), we can obtain the equation of motion for ${\gamma}_{\lambda}(\tau,\mathbf{k})$ as
$u_{k}^{\prime\prime}+\left(k^{2}-\frac{a^{\prime\prime}}{a}\right)u_{k}=0\,,$ (10)
where $u_{k}\equiv\gamma_{\lambda}(\tau,k)aM_{p}/2$ and a prime denotes $d/d\tau$. The power spectrum of primordial GWs is defined as
$P_{T}={k^{3}\over 2\pi^{2}}\sum_{\lambda=+,\times}|\gamma_{\lambda}|^{2}=\frac{4k^{3}}{\pi^{2}M_{p}^{2}}\cdot\frac{\left|u_{k}\right|^{2}}{a^{2}}\,,$ (11)
which should be evaluated at the end of the second stage of slow-roll inflation.
We numerically solve Eq. (10) and obtain the power spectrum of primordial GWs for our model. The results are displayed in Fig. 6. We observe a significant enhancement in the power spectrum amplitude of perturbation modes that exit the horizon around the NEC violation epoch in this model. This is because perturbation modes that are outside the horizon during the NEC violation period experience rapid growth, surpassing the amplitude of the constant mode. Therefore, the power spectrum at intermediate scales exhibits a blue tilt. Similar to the model in Ref. Cai:2020qpu , the model in this paper also has the potential to explain observations at the PTA scale. However, unlike the model in Ref. Cai:2020qpu , the power spectrum amplitudes of perturbation modes that exit the horizon during the second stage of slow-roll inflation are not significantly different from those that exit during the first stage.
Figure 6: The power spectrum $P_{T}$ of primordial GWs for our model obtained by numerically solving Eq. (10). We have set $q_{1}=10$, $q_{2}=6$, $q_{3}=10$, $q_{4}=4$, $\phi_{0}=3.2$, $\phi_{1}=2$, $\phi_{2}=-4$, $\phi_{3}=-4.38$, $\phi_{4}=20$, $\phi_{5}=29$, $f_{1}=1$, $f_{2}=40$, $\lambda=0.01$, $\sigma=23$, $m=-4.5\times 10^{-6}$, $\phi(t_{\rm ini})=-6$, $\dot{\phi}(t_{\rm ini})=0$ and $t_{\rm ini}=0$.
In the following, we attempt to provide an analytical derivation of the power spectrum. Analogous to the approach in Ref. Cai:2020qpu , we divide the background evolution in this model into four stages, with the scale factor of the $j$-th stage approximately parameterized as
$a_{j}(\tau)\sim\left({\tau}_{R,j}-\tau\right)^{\frac{1}{\epsilon_{j}-1}},$ (12)
where $\epsilon_{j}\equiv-\dot{H_{j}}/H_{j}^{2}={3\over 2}(1+w_{j})$, ${\tau}_{R,j}\equiv\tau_{j}-(\epsilon_{j}-1)^{-1}\mathcal{H}^{-1}(\tau_{j})$, $j=1$, 2, 3, 4. For simplicity, we assume $\epsilon_{j}$ (i.e., the equation of state parameter $w_{j}$) as a constant. We also require continuity of $a$ and $a^{\prime}$ at each transition moment.
With Eq. (12), we have
$\frac{a_{j}^{\prime\prime}}{a_{j}}={\nu_{j}^{2}-{1/4}\over{(\tau-\tau_{R,j})^{2}}}\,,$ (13)
where $\nu_{j}={3\over 2}\left|{1-w_{j}\over 1+3w_{j}}\right|$. Consequently, the solutions to Eq. (10) for each phase can be given as
$u_{k,1}(\tau)$ $=$ ${\sqrt{\pi(\tau_{R,1}-\tau)}\over 2}\left\{\alpha_{1}H_{\nu_{1}}^{(1)}[k(\tau_{R,1}-\tau)]+\beta_{1}H_{\nu_{1}}^{(2)}[k(\tau_{R,1}-\tau)]\right\},\,\,(\tau<\tau_{1})\,,$ (14)
$u_{k,{2}}(\tau)$ $=$ ${\sqrt{\pi(\tau_{R,{2}}-\tau)}\over 2}\Big{\{}\alpha_{2}H_{\nu_{2}}^{(1)}[k(\tau_{R,2}-\tau)]+\beta_{2}H_{\nu_{2}}^{(2)}[k(\tau_{R,2}-\tau)]\Big{\}},(\tau_{1}\leq\tau\leq\tau_{2})\,,$
$u_{k,{3}}(\tau)$ $=$ ${\sqrt{\pi(\tau_{R,{3}}-\tau)}\over 2}\Big{\{}\alpha_{3}H_{\nu_{3}}^{(1)}[k(\tau_{R,3}-\tau)]+\beta_{3}H_{\nu_{3}}^{(2)}[k(\tau_{R,3}-\tau)]\Big{\}},(\tau_{2}<\tau\leq\tau_{3})\,,$
$u_{k,{4}}(\tau)$ $=$ ${\sqrt{\pi(\tau_{R,{4}}-\tau)}\over 2}\Big{\{}\alpha_{4}H_{\nu_{4}}^{(1)}[k(\tau_{R,4}-\tau)]+\beta_{4}H_{\nu_{4}}^{(2)}[k(\tau_{R,4}-\tau)]\Big{\}},(\tau_{3}<\tau\leq\tau_{4})\,,$
where $\alpha_{j}$ and $\beta_{j}$ are $k$-dependent coefficients, $j=1$, 2, 3, 4. We set the initial state as the Bunch-Davis vacuum, i.e., $u_{k}=e^{-ik\tau}/\sqrt{2k}$, which indicates that $|\alpha_{1}|=1$ and $|\beta_{1}|=0$.
Using the matching conditions $u_{k,j}(\tau_{j})=u_{k,j+1}(\tau_{j})$ and $u_{k,j}^{\prime}\big{|}{\tau=\tau{j}}=u_{k,j+1}^{\prime}\big{|}{\tau=\tau{j}}$, we can obtain all of the $\alpha_{j}$ and $\beta_{j}$ for $j=2$, 3, 4. We have
$\left(\begin{array}[]{ccc}\alpha_{j+1}\\ \beta_{j+1}\end{array}\right)$ $=$ ${\cal M}^{(j)}\left(\begin{array}[]{ccc}\alpha_{j}\\ \beta_{j}\end{array}\right)\,,\quad{\rm where}\quad{\cal M}^{(j)}=\left(\begin{array}[]{ccc}{\cal M}^{(j)}_{11}&{\cal M}^{(j)}_{12}\\ {\cal M}^{(j)}_{21}&{\cal M}^{(j)}_{22}\end{array}\right)\,,$ (21)
The details of the matrix elements of the transfer matrix can be found in Cai:2015nya ; Cai:2019hge . At the end of the second slow-roll inflation, the power spectrum can be given as
$P_{T}=\frac{4k^{3}}{\pi^{2}M_{p}^{2}}\cdot\frac{\left|u_{k,4}\right|^{2}}{a^{2}}\Big{|}_{\tau\rightarrow\tau_{4}}=P_{T,inf2}\left|\alpha_{4}-\beta_{4}\right|^{2}\,,$ (22)
where $P_{T,inf2}={2H_{inf2}^{2}\over\pi^{2}M_{p}^{2}}$. A detailed expression for $\alpha_{4}$ and $\beta_{4}$ can be found in Appendix A. We plot $P_{T}/P_{T,\text{inf2}}=\left|\alpha_{4}-\beta_{4}\right|^{2}$ in Fig. 7 for different choices of $w_{i}$. We can see from Figs. 6 and 7(a) that the approximation works well for our model with appropriate choices of $w_{i}$. In our parameterization, the height difference between the two platforms in the potential can be adjusted by changing $\epsilon_{3}=-\dot{H}/H^{2}={3\over 2}(1+w_{j})$.
Figure 7: The power spectrum of the primordial GWs given by Eq. (22). For the left panel, we set $w_{1}=-1$, $w_{2}=-10$, $w_{3}=1$, and $w_{4}=-1$, providing a close approximation to the result shown in Fig. 6. In the right panel, we choose $w_{1}=-1$, $w_{2}=-8$, $w_{3}=2/3$, and $w_{4}=-1$, resembling the scenario discussed in Ref. Cai:2020qpu .
IV The energy density spectrum of GWs
The energy density spectrum of GWs Turner:1993vb ; Boyle:2005se (see also Zhao:2006mm ; Kuroyanagi:2014nba ) is
$\Omega_{\rm{GW}}(\tau_{0})=\frac{k^{2}}{12a_{0}^{2}H^{2}_{0}}P_{T}(k)\left[\frac{3\Omega_{{m}}j_{1}(k\tau_{0})}{k\tau_{0}}\sqrt{1.0+1.36\frac{k}{k_{\text{eq}}}+2.50\left(\frac{k}{k_{\text{eq}}}\right)^{2}}\right]^{2},$ (23)
where $H_{0}=67.8\,{\rm km/s/Mpc}$, $\tau_{0}=1.41\times 10^{4}$ Mpc, $a_{0}=1$, $k_{\text{eq}}=0.073\,\Omega_{\text{m}}h^{2}$ Mpc-1 is the wavenumber corresponding to the mode that entered the horizon at the equality of matter and radiation, $\Omega_{\rm m}=\rho_{m}/\rho_{c}$ is the density fraction of matter today, $\rho_{{c}}=3H^{2}_{0}/\big{(}8\pi G\big{)}$ is the critical energy density, and $j_{1}(x)$ is the spherical Bessel function of the first kind. We plot the energy density spectrum of GWs for our model in Fig. 8, using the numerical results obtained by solving Eq. (10). Namely, Fig. 8 correspond to the power spectrum displayed in Fig. 6.
Figure 8: The energy density spectrum of GWs for our model. The blue curve is obtained by using the numerical results obtained by solving Eq. (10), while the red curve is based on the power spectrum shown in Fig. 7(b). We have set $P_{T,inf2}=1.82\times 10^{-10}$.
For observational purposes, the distinctive features may include the blue-tilted nature in the middle-frequency range and the decreasing amplitude of these oscillations at higher frequencies. Recently, observational data reported by NANOGrav NANOGrav:2023gor ; NANOGrav:2023hvm ; NANOGrav:2020bcs , EPTA EPTA:2023fyk , PPTA Reardon:2023gzh and CPTA Xu:2023wog is consistent with a stochastic GW background with an amplitude of $\Omega_{\rm{GW}}\simeq 10^{-9}$ and a spectral index of $n_{T}=1.8\pm 0.3$ at the reference frequency $f=1$ $\text{yr}^{-1}$. If the observational data are interpreted as primordial GW background, it corresponds to a strongly blue-tilted spectrum with an amplitude of $P_{T}\simeq 10^{6}\Omega_{\rm{GW}}\simeq 10^{-3}$ at a frequency of $f=1$ $\text{yr}^{-1}$. We have plotted the energy density spectrum in red in Fig. 8 using the analytical formula given in Eq. (22) for a specific set of parameters. Fig. 8 suggests that future PTA observations may have the potential to detect or constrain our model through the specific features of the power spectrum, which requires further investigation.
V Conclusion
In this paper, we propose a new single-field model where the universe sequentially undergoes a first stage of slow-roll inflation, NEC violation, and a second stage of slow-roll inflation. Within the potential responsible for slow-roll inflation, there exists a high barrier, over which the rolling inflaton climb via the NEC violation mechanism [^5] See also Caravano:2024tlp for other mechanism. before entering the second inflationary stage. Following the NEC violation, there is a process of energy scale reduction, resulting in the energy scales of the two slow-roll inflation stages being relatively close, contrasting with the model in Cai:2020qpu .
We calculate the power spectrum of primordial GWs using both numerical and analytical methods, comparing the results obtained from these two approaches. We find that the approximation used in the analytical derivation closely matches the results from numerical calculations. The power spectrum is nearly scale-invariant on both large and small scales. However, in the middle-frequency range, it shows a blue-tilted nature, while at higher frequencies, there is a reduction in oscillation amplitudes. These unique characteristics could potentially differentiate our model from others.
In the early universe, NEC violation may have occurred multiple times. We present a sketch of the GW power spectrum in Fig. 9 for the scenario where $\phi$ climbs over the potential barriers multiple times. The accumulation of future observational data from various frequency bands, including CMB, PTA, Advanced LIGO, Advanced Virgo, LISA, Taiji and Tianqin, will enable more tests or constraints on NEC violation in the very early universe.
Figure 9: A sketch of the GW power spectrum for the scenario where $\phi$ climbs over the potential barriers multiple times.
Acknowledgements.
Y. C. is supported in part by the Natural Science Foundation of Henan Province and Zhengzhou University (Grant Nos. 242300420231, JC23149007, 35220136), the China Postdoctoral Science Foundation (Grant No. 2021M692942). Y.-S. P. is supported by NSFC, No.12075246, National Key Research and Development Program of China, No. 2021YFC2203004, and the Fundamental Research Funds for the Central Universities.
Appendix A Detailed analytic expression of $\alpha$ and $\beta$
The scale factor can be parameterized as a piecewise function of $\tau$, i.e.,
$a_{j}(\tau)=a_{j}(\tau_{j})\left(\frac{\tau-\tau_{R,j}}{\tau_{j}-\tau_{R,j}}\right)^{\frac{1}{\epsilon_{j}-1}}~{},\quad\tau<\tau_{j}\,.$ (24)
The continuity of $a$ gives
$a_{1}(\tau_{1})=a_{2}(\tau_{2})\left(\frac{\tau_{1}-\tau_{R,2}}{\tau_{2}-\tau_{R,2}}\right)^{\frac{1}{\epsilon_{2}-1}},$ (25)
$a_{2}(\tau_{2})=a_{3}(\tau_{3})\left(\frac{\tau_{2}-\tau_{R,3}}{\tau_{3}-\tau_{R,3}}\right)^{\frac{1}{\epsilon_{3}-1}},~{}$ (26)
$a_{3}(\tau_{3})=a_{4}(\tau_{4})\left(\frac{\tau_{3}-\tau_{R,4}}{\tau_{4}-\tau_{R,4}}\right)^{\frac{1}{\epsilon_{4}-1}}.$ (27)
Considering the continuity of $a^{\prime}$ or $\mathcal{H}$, we can define the following quantities:
¯$\bar{\mathcal{H}}_{1}\equiv\mathcal{H}(\tau_{1})~{},~{}\bar{\mathcal{H}}_{2}\equiv\mathcal{H}(\tau_{2})~{},~{}\bar{\mathcal{H}}_{3}\equiv\mathcal{H}(\tau_{3})~{}.$ (28)
Using (24), we can obtain the integration constants $\tau_{R,j}$ as
$\tau_{R,1}=\tau_{1}+\bar{\mathcal{H}}_{1}^{-1}~{},~{}\tau_{R,2}=\tau_{1}+({v_{2}-\frac{1}{2}}){\bar{\mathcal{H}}_{1}^{-1}},\tau_{R,2}=\tau_{2}+({v_{2}-\frac{1}{2}}){\bar{\mathcal{H}}_{2}^{-1}},$ (29)
$\tau_{R,3}=\tau_{2}+({-v_{3}-\frac{1}{2}}){\bar{\mathcal{H}}_{2}^{-1}},\tau_{R,3}=\tau_{3}+({-v_{3}-\frac{1}{2}}){\bar{\mathcal{H}}_{3}^{-1}},~{}\tau_{R,4}=\tau_{3}+\bar{\mathcal{H}}_{3}^{-1},$ (30)
where $\nu_{j}={3\over 2}\left|{1-w_{j}\over 1+3w_{j}}\right|$, and $\epsilon_{j}={3\over 2}(1+w_{j})\simeq{\rm const.}$. Note that for $-{1\over 3}<w_{j}<1$ and $w_{j}<-{1\over 3}$, we have $\nu_{j}=-{1\over 2}-{1\over 1-\epsilon_{j}}$ and $\nu_{j}={1\over 2}+{1\over 1-\epsilon_{j}}$, respectively.
Using the matching method, we obtain the following recursive relations:
$~{}~{}~{}~{}~{}~{}\alpha_{2}$ $=\frac{i\pi}{8x_{1}}\sqrt{\frac{x_{1}}{x_{2}}}\left(H_{\frac{3}{2}}^{(1)}(x_{1})\left(2x_{1}x_{2}H_{\nu_{2}-1}^{(2)}(x_{2})+\left(-2\nu_{2}x_{1}+x_{1}+2x_{2}\right)H_{\nu_{2}}^{(2)}(x_{2})\right)\right.$
$\quad-2x_{1}x_{2}H_{\frac{1}{2}}^{(1)}(x_{1})H_{\nu_{2}}^{(2)}(x_{2})\Big{)},$ (31)
$~{}~{}~{}~{}~{}~{}~{}~{}~{}\beta_{2}$ $=-\frac{i\pi}{8x_{1}}\sqrt{\frac{x_{1}}{x_{2}}}\left(H_{\frac{3}{2}}^{(1)}(x_{1})\left(2x_{1}x_{2}H_{\nu_{2}-1}^{(1)}(x_{2})+\left(-2\nu_{2}x_{1}+x_{1}+2x_{2}\right)H_{\nu_{2}}^{(1)}(x_{2})\right)\right.$
$\quad-2x_{1}x_{2}H_{\frac{1}{2}}^{(1)}(x_{1})H_{\nu_{2}}^{(1)}(x_{2})\Big{)},$ (32)
$\alpha_{3}$ $=-\frac{\pi}{8y_{1}}\sqrt{-\frac{y_{1}}{y_{2}}}\Big{(}y_{2}H_{\nu_{3}}^{(2)}(y_{2})\Big{(}2\alpha_{2}y_{1}H_{\nu_{2}-1}^{(1)}(y_{1})+(\alpha_{2}-2\alpha_{2}\nu_{2})H_{\nu_{2}}^{(1)}(y_{1})$
$\quad+2\beta_{2}y_{1}H_{\nu_{2}-1}^{(2)}(y_{1})+(\beta_{2}-2\beta_{2}\nu_{2})H_{\nu_{2}}^{(2)}(y_{1})\Big{)}$
$\quad-y_{1}\Big{(}2y_{2}H_{\nu_{3}-1}^{(2)}(y_{2})+(1-2\nu_{3})H_{\nu_{3}}^{(2)}(y_{2})\Big{)}$
$\quad\cdot\Big{(}\alpha_{2}H_{\nu_{2}}^{(1)}(y_{1})+\beta_{2}H_{\nu_{2}}^{(2)}(y_{1})\Big{)}\Big{)},$ (33)
$\beta_{3}$ $=\frac{\pi}{8y_{1}}\sqrt{-\frac{y_{1}}{y_{2}}}\Big{(}2\alpha_{2}y_{1}y_{2}H_{\nu_{2}-1}^{(1)}(y_{1})H_{\nu_{3}}^{(1)}(y_{2})$
$\quad+\alpha_{2}H_{\nu_{2}}^{(1)}(y_{1})\Big{(}\big{(}2\nu_{3}y_{1}-2\nu_{2}y_{2}-y_{1}+y_{2}\big{)}H_{\nu_{3}}^{(1)}(y_{2})-2y_{1}y_{2}H_{\nu_{3}-1}^{(1)}(y_{2})\Big{)}$
$\quad-2\beta_{2}y_{1}y_{2}H_{\nu_{3}-1}^{(1)}(y_{2})H_{\nu_{2}}^{(2)}(y_{1})+\beta_{2}H_{\nu_{3}}^{(1)}(y_{2})\Big{(}2y_{1}y_{2}H_{\nu_{2}-1}^{(2)}(y_{1})$
$\quad+\big{(}2\nu_{3}y_{1}-2\nu_{2}y_{2}-y_{1}+y_{2}\big{)}H_{\nu_{2}}^{(2)}(y_{1})\Big{)}\Big{)},$ (34)
$\alpha_{4}$ $=\frac{\pi}{8z_{1}}\sqrt{-\frac{z_{1}}{z_{2}}}\Big{(}2z_{1}\Big{(}H_{\frac{3}{2}}^{(2)}(z_{2})-z_{2}H_{\frac{1}{2}}^{(2)}(z_{2})\Big{)}\Big{(}\alpha_{3}H_{\nu_{3}}^{(1)}(z_{1})+\beta_{3}H_{\nu_{3}}^{(2)}(z_{1})\Big{)}$
$\quad+z_{2}H_{\frac{3}{2}}^{(2)}(z_{2})\Big{(}2\alpha_{3}z_{1}H_{\nu_{3}-1}^{(1)}(z_{1})+\big{(}\alpha_{3}-2\alpha_{3}\nu_{3}\big{)}H_{\nu_{3}}^{(1)}(z_{1})$
$\quad+2\beta_{3}z_{1}H_{\nu_{3}-1}^{(2)}(z_{1})+\big{(}\beta_{3}-2\beta_{3}\nu_{3}\big{)}H_{\nu_{3}}^{(2)}(z_{1})\Big{)}\Big{)},$ (35)
$\begin{split}\beta_{4}&=-\frac{\pi}{8z_{1}}\sqrt{-\frac{z_{1}}{z_{2}}}\Big{(}H_{\frac{3}{2}}^{(1)}(z_{2})\Big{(}2\alpha_{3}z_{1}z_{2}H_{\nu_{3}-1}^{(1)}(z_{1})\\ &\quad+\alpha_{3}\big{(}-2\nu_{3}z_{2}+2z_{1}+z_{2}\big{)}H_{\nu_{3}}^{(1)}(z_{1})\\ &\quad+2\beta_{3}z_{1}z_{2}H_{\nu_{3}-1}^{(2)}(z_{1})+\beta_{3}\big{(}-2\nu_{3}z_{2}+2z_{1}+z_{2}\big{)}H_{\nu_{3}}^{(2)}(z_{1})\Big{)}\\ &\quad-2z_{1}z_{2}H_{\frac{1}{2}}^{(1)}(z_{2})\Big{(}\alpha_{3}H_{\nu_{3}}^{(1)}(z_{1})+\beta_{3}H_{\nu_{3}}^{(2)}(z_{1})\Big{)}\Big{)},\end{split}$ (36)
where $x_{1}=k(\tau_{R,1}-\tau_{1})$, $x_{2}=k(\tau_{R,2}-\tau_{1})$, $y_{1}=k(\tau_{R,2}-\tau_{2})$, $y_{2}=k(\tau_{R,3}-\tau_{2})$, $z_{1}=k(\tau_{R,3}-\tau_{3})$, $z_{2}=k(\tau_{R,4}-\tau_{3})$. Since the results are too complex, we only provide the recursive formulas for $\alpha_{j}$ and $\beta_{j}$.
Figure 10: Left: Comparison between $|G_{1}|^{2}$ (green dashed curve) and the numerical result of ${P_{T}/P_{T,inf2}}$ (blue dotted curve). Right: Comparison between the template (38) (green dashed curve) and the numerical result of ${P_{T}/P_{T,inf2}}$ (blue dotted curve). We have set ¯$\bar{\cal H}_{1}=10^{4}a_{0}H_{0}$, where $a_{0}=1$ and $H_{0}=67\,{\rm km/s/Mpc}$. We also set $w_{2}=-15$ and $w_{3}={2}/{3}$, which corresponding to $\nu_{2}={6}/{11}$ and $\nu_{3}={1}/{6}$, respectively.
Next, we attempt to explore potential approximations. It’s worth noting that in our model, $y_{2}$ and $z_{1}$ are negative. Therefore, our approach here is to consider $k\gg\bar{\mathcal{H}}_{1}$. By series expansion, we get
$G_{1}$ $={\alpha_{4}-\beta_{4}}$ (37)
$=\frac{\pi^{3/2}}{16\sqrt{2}z_{2}^{2}}\bigg{(}2H_{\nu_{2}-1}^{(1)}\left(y_{1}\right)\bigg{(}H_{\nu_{3}}^{(1)}\left(y_{2}\right)\bigg{(}2H_{\nu_{3}-1}^{(2)}\left(z_{1}\right)\sqrt{y_{1}y_{2}z_{1}}z_{2}\left(\cos\left(z_{2}\right)z_{2}-\sin\left(z_{2}\right)\right)$
$+H_{\nu_{3}}^{(2)}\left(z_{1}\right)\bigg{(}\sin\left(z_{2}\right)\left(2\sqrt{y_{1}y_{2}z_{1}}z_{2}^{2}+\sqrt{\frac{y_{1}y_{2}}{z_{1}}}\left(1-2\nu_{3}\right)z_{2}-2\sqrt{y_{1}y_{2}z_{1}}\right)$
$+\cos\left(z_{2}\right)z_{2}\left(\sqrt{\frac{y_{1}y_{2}}{z_{1}}}z_{2}\left(2\nu_{3}-1\right)+2\sqrt{y_{1}y_{2}z_{1}}\right)\bigg{)}\bigg{)}$
$+H_{\nu_{3}}^{(2)}\left(y_{2}\right)\bigg{(}H_{\nu_{3}}^{(1)}\left(z_{1}\right)\bigg{(}\sin\left(z_{2}\right)\left(-2\sqrt{y_{1}y_{2}z_{1}}z_{2}^{2}+\sqrt{\frac{y_{1}y_{2}}{z_{1}}}\left(2\nu_{3}-1\right)z_{2}+2\sqrt{y_{1}y_{2}z_{1}}\right)$
$-\cos\left(z_{2}\right)z_{2}\left(\sqrt{\frac{y_{1}y_{2}}{z_{1}}}z_{2}\left(2\nu_{3}-1\right)+2\sqrt{y_{1}y_{2}z_{1}}\right)\bigg{)}-2H_{\nu_{3}-1}^{(1)}\left(z_{1}\right)\sqrt{y_{1}y_{2}z_{1}}z_{2}\big{(}\cos\left(z_{2}\right)z_{2}$
$-\sin\left(z_{2}\right)\big{)}\bigg{)}\bigg{)}+H_{\nu_{2}}^{(1)}\left(y_{1}\right)\bigg{(}-4\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{1}}{y_{2}z_{1}}}\nu_{3}^{2}z_{2}^{2}$
$+2\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{2}}{y_{1}z_{1}}}\nu_{2}z_{2}^{2}-4\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}-1}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{2}z_{1}}{y_{1}}}\nu_{2}z_{2}^{2}$
$-4H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{2}z_{1}}{y_{1}}}\nu_{2}z_{2}^{2}-4\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{2}}{y_{1}z_{1}}}\nu_{2}\nu_{3}z_{2}^{2}$
$+4\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{1}}{y_{2}z_{1}}}\nu_{3}z_{2}^{2}+2\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{2}}{y_{1}z_{1}}}\nu_{3}z_{2}^{2}$
$-4\cos\left(z_{2}\right)H_{\nu_{3}-1}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{1}y_{2}}{z_{1}}}\nu_{3}z_{2}^{2}-4\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}-1}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{1}z_{1}}{y_{2}}}\nu_{3}z_{2}^{2}$
$-4H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{1}z_{1}}{y_{2}}}\nu_{3}z_{2}^{2}-\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{1}}{y_{2}z_{1}}}z_{2}^{2}$
$-\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{2}}{y_{1}z_{1}}}z_{2}^{2}+2\cos\left(z_{2}\right)H_{\nu_{3}-1}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{1}y_{2}}{z_{1}}}z_{2}^{2}$
$+2\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}-1}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{1}z_{1}}{y_{2}}}z_{2}^{2}+2H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{1}z_{1}}{y_{2}}}z_{2}^{2}$
$+2\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}-1}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{2}z_{1}}{y_{1}}}z_{2}^{2}+2H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{2}z_{1}}{y_{1}}}z_{2}^{2}$
$-4\cos\left(z_{2}\right)H_{\nu_{3}-1}^{(1)}\left(y_{2}\right)H_{\nu_{3}-1}^{(2)}\left(z_{1}\right)\sqrt{y_{1}y_{2}z_{1}}z_{2}^{2}-4H_{\nu_{3}-1}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{y_{1}y_{2}z_{1}}z_{2}^{2}$
$+4H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{1}}{y_{2}z_{1}}}\nu_{3}^{2}z_{2}-2H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{2}}{y_{1}z_{1}}}\nu_{2}z_{2}$
$-4\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{2}z_{1}}{y_{1}}}\nu_{2}z_{2}+4H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}-1}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{2}z_{1}}{y_{1}}}\nu_{2}z_{2}$
$+4H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{2}}{y_{1}z_{1}}}\nu_{2}\nu_{3}z_{2}-4H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{1}}{y_{2}z_{1}}}\nu_{3}z_{2}$
$-2H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{2}}{y_{1}z_{1}}}\nu_{3}z_{2}+4H_{\nu_{3}-1}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{1}y_{2}}{z_{1}}}\nu_{3}z_{2}$
$\begin{split}&-4\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{1}z_{1}}{y_{2}}}\nu_{3}z_{2}+4H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}-1}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{1}z_{1}}{y_{2}}}\nu_{3}z_{2}\\ &+2H_{\nu_{3}-1}^{(1)}\left(z_{1}\right)\left(\cos\left(z_{2}\right)z_{2}-\sin\left(z_{2}\right)\right)\bigg{(}2\sqrt{y_{1}y_{2}z_{1}}H_{\nu_{3}-1}^{(2)}\left(y_{2}\right)+H_{\nu_{3}}^{(2)}\left(y_{2}\right)\bigg{(}2\sqrt{\frac{y_{2}z_{1}}{y_{1}}}\nu_{2}\\ &+2\sqrt{\frac{y_{1}z_{1}}{y_{2}}}\nu_{3}-\sqrt{\frac{y_{1}z_{1}}{y_{2}}}-\sqrt{\frac{y_{2}z_{1}}{y_{1}}}\bigg{)}\bigg{)}z_{2}+H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{1}}{y_{2}z_{1}}}z_{2}\\ &+H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{2}}{y_{1}z_{1}}}z_{2}-2H_{\nu_{3}-1}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{1}y_{2}}{z_{1}}}z_{2}\\ &+2\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{1}z_{1}}{y_{2}}}z_{2}-2H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}-1}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{1}z_{1}}{y_{2}}}z_{2}\\ &+2\cos\left(z_{2}\right)H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{\frac{y_{2}z_{1}}{y_{1}}}z_{2}-2H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}-1}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{2}z_{1}}{y_{1}}}z_{2}\\ &-4\cos\left(z_{2}\right)H_{\nu_{3}-1}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sqrt{y_{1}y_{2}z_{1}}z_{2}+4H_{\nu_{3}-1}^{(1)}\left(y_{2}\right)H_{\nu_{3}-1}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{y_{1}y_{2}z_{1}}z_{2}\\ &+4H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{2}z_{1}}{y_{1}}}\nu_{2}+4H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{1}z_{1}}{y_{2}}}\nu_{3}\\ &+H_{\nu_{3}}^{(1)}\left(z_{1}\right)\bigg{(}2H_{\nu_{3}-1}^{(2)}\left(y_{2}\right)\bigg{(}\sin\left(z_{2}\right)\bigg{(}2\sqrt{y_{1}y_{2}z_{1}}z_{2}^{2}+\sqrt{\frac{y_{1}y_{2}}{z_{1}}}\left(1-2\nu_{3}\right)z_{2}-2\sqrt{y_{1}y_{2}z_{1}}\bigg{)}\\ &+\cos\left(z_{2}\right)z_{2}\left(\sqrt{\frac{y_{1}y_{2}}{z_{1}}}z_{2}\left(2\nu_{3}-1\right)+2\sqrt{y_{1}y_{2}z_{1}}\right)\bigg{)}+H_{\nu_{3}}^{(2)}\left(y_{2}\right)\bigg{(}\cos\left(z_{2}\right)z_{2}\bigg{(}4\sqrt{\frac{y_{2}z_{1}}{y_{1}}}\nu_{2}\\ &+4\sqrt{\frac{y_{1}z_{1}}{y_{2}}}\nu_{3}+z_{2}\left(2\nu_{3}-1\right)\left(2\sqrt{\frac{y_{2}}{y_{1}z_{1}}}\nu_{2}+2\sqrt{\frac{y_{1}}{y_{2}z_{1}}}\nu_{3}-\sqrt{\frac{y_{1}}{y_{2}z_{1}}}-\sqrt{\frac{y_{2}}{y_{1}z_{1}}}\right)-2\sqrt{\frac{y_{1}z_{1}}{y_{2}}}\\ &-2\sqrt{\frac{y_{2}z_{1}}{y_{1}}}\bigg{)}+\sin\left(z_{2}\right)\bigg{(}\bigg{(}-2\bigg{(}\sqrt{\frac{y_{1}z_{1}}{y_{2}}}+\sqrt{\frac{y_{2}z_{1}}{y_{1}}}\bigg{)}+4\sqrt{\frac{y_{2}z_{1}}{y_{1}}}\nu_{2}+4\sqrt{\frac{y_{1}z_{1}}{y_{2}}}\nu_{3}\bigg{)}z_{2}^{2}\\ &-\left(2\nu_{3}-1\right)\left(2\sqrt{\frac{y_{2}}{y_{1}z_{1}}}\nu_{2}+2\sqrt{\frac{y_{1}}{y_{2}z_{1}}}\nu_{3}-\sqrt{\frac{y_{1}}{y_{2}z_{1}}}-\sqrt{\frac{y_{2}}{y_{1}z_{1}}}\right)z_{2}+2\bigg{(}-2\sqrt{\frac{y_{2}z_{1}}{y_{1}}}\nu_{2}\\ &-2\sqrt{\frac{y_{1}z_{1}}{y_{2}}}\nu_{3}+\sqrt{\frac{y_{1}z_{1}}{y_{2}}}+\sqrt{\frac{y_{2}z_{1}}{y_{1}}}\bigg{)}\bigg{)}\bigg{)}\bigg{)}-2H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{1}z_{1}}{y_{2}}}\\ &-2H_{\nu_{3}}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{\frac{y_{2}z_{1}}{y_{1}}}+4H_{\nu_{3}-1}^{(1)}\left(y_{2}\right)H_{\nu_{3}}^{(2)}\left(z_{1}\right)\sin\left(z_{2}\right)\sqrt{y_{1}y_{2}z_{1}}\bigg{)}\bigg{)}.\end{split}$
This expression remains quite complex, although it is much simpler compared to the original expression for $\alpha_{4}-\beta_{4}$. We plot $|\alpha_{4}-\beta_{4}|^{2}=|G_{1}|^{2}$ in Fig. 10(a). It can be observed that, in the limit where $k\gg\bar{\mathcal{H}}_{1}$, this approximation matches the numerical solution quite well.
To better match the numerical results across the entire frequency range, we propose a parameterized template as
${P_{T}(k)\over P_{T,inf2}}=\frac{1}{1+e^{k-{\bar{\mathcal{H}}}_{3}}}+|G_{1}|^{2}\,.$ (38)
We compare this template with the numerical results in Fig. 10(b). It is evident that Eq. (38) serves as a reliable approximation for observational purposes.